# Blind Image Deblurring using FFT-ReLU with Deep Learning Pipeline Integration

### v1

@@ Abstract
Blind image deblurring is the process of deriving a sharp image and a blur kernel from a blurred image. Blurry images are typically modeled as the convolution of a sharp image with a blur kernel, necessitating the estimation of the unknown blur kernel to perform blind image deblurring effectively. Existing approaches primarily focus on domain-specific features of images, such as salient edges, dark channels, and light streaks. These features serve as probabilistic priors to enhance the estimation of the blur kernel. For improved generality, we propose a novel prior (ReLU sparsity prior) that estimates blur kernel effectively across all distributions of images (natural, facial, text, low-light, saturated etc). Our approach demonstrates superior efficiency, with inference times up to three times faster, while maintaining high accuracy in PSNR, SSIM, and error ratio metrics. We also observe noticeable improvement in the performance of the state-of-the-art architectures (in terms of aforementioned metrics) in deep learning based approaches when our method is used as a post-processing unit.
keywords:
Image restoration, Fast Fourier Transform, Kernel estimation, Blind deconvolution
@@ 1 Introduction
Figure 1: Results of our blind image deblurring algorithm on two blurry images
Blind image deblurring is a classical problem in the fields of image and signal processing, and has been a subject of active research in the fields of computer vision. With more and more handheld devices being used to capture photos every day, blurring has become a significant source of degradation of image quality. The types of blur encountered include Gaussian blur, motion blur, out-of-focus blur, atmospheric blur, and others. Motion blur and out-of-focus blur are identified as the primary sources of blur in image restoration.
The objective of blind image deblurring is to estimate a blur kernel and a latent sharp image from the input blurry image. In the case of spatially invariant and uniform blur, the process of blurring can be expressed as follows:
$B=I\otimes K+n$
where $B$ is the input blurry image, $I$ is the latent sharp image, $\otimes$ denotes convolution, $K$ is the blur kernel i.e. the point spread function (PSF) controlling the degree of distortion of the image, and $n$ is the additional noise that is gathered during image acquisition. Since the only available input is $B$, the problem becomes ill-posed: many pairs of $I$ and $K$ can produce the same $B$.
A common approach taken in blind image deblurring is the maximum a posteriori (MAP) framework, where the latent image and the blur kernel are optimized simultaneously in turns. Different assumptions or heuristics involving the latent image, the kernel, or both are used in such frameworks from [4, 11, 17, 22, 31, 36, 49], designing new priors that provide effective results. However, these methods are generally handcrafted for images from individual domains, such as the work of [31], since they build on image-specific properties that aid in optimization techniques. Moreover, the calculation of some of these priors can incur high computational costs, as noted by [32], which may reduce the effectiveness in deblurring images. Since the blur kernel and the latent image are typically solved in an alternating fashion, the initialization steps play a vital role in the convergence of any deblurring algorithm. In the frequency domain, the blurred image is instantiated as the multiplications of the Fourier transforms of the latent image and the blur kernel (PSF), added with sensor noise, as proposed in [3, 49, 14, 32].
$B(\omega)=I(\omega)K(\omega)+N(\omega)$
where $B(\omega)$, $I(\omega)$, $K(\omega)$ and $N(\omega)$ represent Fourier transforms of the blurry image, the latent sharp image, the blur kernel, and the sensor noise, respectively, and $\omega\in[-\pi,\pi]^{2}$. [24] depicts that a series of Fourier Transform, ReLU, and Inverse Fourier Transform operations can implicitly reveal faithful information about the blur kernel, such as the direction and the level of the blur. They proposed a Res FFT-ReLU block for performing these operations, which can be placed and used in existing architectures based on neural networks. However, its applications in blind image deblurring remained unexplored.
In this paper, we propose a new prior, the ReLU Sparsity prior. A sharp image has a higher number of zero elements compared to its blurry counterpart, which is a property that can lead to producing the latent image, as shown empirically in our work. Besides, we refer to the aforementioned sequence of operations together as RFT and use its L0 norm in our prior. Minimizing the objective function would necessarily imply penalizing the lower sparsity of the blurry image. Since an L0 norm is involved, we use the half quadratic splitting technique, [47], to optimize our objective function.
Another unexplored field of research within the domain of Image Deblurring, in general, is the combination of traditional blind deconvolution methods with deep learning architectures. This gap exists because no previous work has implemented a blind deconvolution method that can be seamlessly applied to transformer-based deep learning architectures. We have mitigated this issue by implementing our method in a pipeline that is suitable for incorporating deep learning.
@@
    We propose a new prior, the ReLU Sparsity prior, which allows us to obtain faithful information about the blur kernel regardless of the nature of the image, while simultaneously applying optimization techniques based on the change in sparsity that occurs due to convolution. Fig. 1 is an example of deblurred output from our algorithm.
@@
    Our algorithm achieves competitive performance based on the PSNR, SSIM and error ratio metrics when compared with state-of-the-art blind image deblurring algorithms from [10, 32, 33, 36, 45, 11, 17], while achieving significantly lower inference time on high-resolution images when compared using images from benchmark datasets prepared by [16, 21].
@@
    We employed our algorithm as a post-processing unit for currently established state-of-the-art methods listed in the literature, including [41, 5, 18, 19, 53, 51, 52], specifically for the RealBlur dataset from [35]. We observed notable improvements in the aforementioned metrics when compared with these works standalone, i.e., without any post-processing units.
2 Related Works
Due to the joint estimation of the blur kernel and the latent sharp image from the input blurry image, single-image deblurring is an under-constrained problem, which is typically formulated in MAP frameworks, such as [27]. Initial advances relied on statistical priors and salient edge detection, such as [10, 45, 49, 36, 11, 17, 22], although the existence of strong edges in latent images is not guaranteed.
Different image priors and likelihood estimations have been proposed to improve the efficiency of MAP frameworks. Such regularizations or assumptions about the blur kernel or the latent image include L0 regularized prior by [31, 48], by dark channel prior [32], extreme channel prior by [50], patch prior by [26], local binary pattern prior by [56], latent structure prior by [2], learned image prior by [23], uniform blur by [21, 49], local maximum difference prior by [25], internal patch recurrence by [28], Laplacian prior by [8], tri-segment intensity prior by [54], graph-based image prior by [1], constant depth by [12, 46], in-plane rotation by [38], and forward motion by [57]. [11] use variational Bayesian inference to learn an image gradient prior. [21] show that this method can avoid trivial solutions that naive MAP-based methods can possibly be unable to.
Sparsity priors have been shown to be useful in kernel estimation in MAP frameworks. [17] used normalized sparsity prior in their MAP framework for kernel estimation, [45] proposed a two-phase technique for deblurring single images: the estimation of blur kernel using edge selections and ISD optimizations is followed by non-blind deconvolution with total variation, with an additional Gaussian prior from [10].
Several newer approaches designed for domain-specific deblurring, such as low-illumination [14], text [7, 9, 31] and face [30] images utilize inherent statistical characteristics of their domains, but display a lack of generality when applied to natural images of other distributions.
Recent advancements in the research of deep learning resulted in the rise of designing neural networks tailored for the task of image restoration, including image deblurring. Besides datasets for blind deblurring algorithms such as [16], datasets were acquired for training and testing these neural networks as well, for example by [37, 29, 35]. Architectures based on CNNs utilised encoder-decoder structures from [40, 52, 15], while GANs made use of their generator-discriminator networks to produce clear images from blurred ones, as proposed by [18, 19, 55]. Later on, architectures based on Transformers such as [42] and Multilayer Perceptrons including [34, 20] also produced excellent results in image restoration.
The advent of deep learning has seen a decrease in new approaches utilising traditional techniques, and as a result, the inherent properties of images and the blurring process is overlooked. The earlier methods, often domain-specific, lack in generalisation; the objective of our research is to design a deblurring framework that combines the strength of these two streams and covers each other’s weakness.
3 ReLU Sparsity Prior
In this section, we discuss the sparsity property of applying RFT operation and develop the ReLU prior to formulate an objective function to estimate the latent sharp image and blur kernel. In Section 3.1, we explain the concept of sparsity and the RFT operations, and the connections between them. We define our model in Section 3.2. In the subsequent sections, we delve into the blind deconvolution and non-blind deconvolution aspects of our algorithm in further detail.
3.1 Sparsity and RFT
To describe our work, we begin by discussing the effect of convolution on the sparsity of images. In standard deblurring models, blurry images are described as the result of convolving a sharp image with a blur kernel (Equation (1)). Here, $(x,y)$ denotes the pixel location, and the blur kernel $K$ has a size of $k\times k$. $S$ and $B$ are the sharp image and the blurry image, respectively.
$B(x,y)=\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}S(x+i,y+j)\cdot K(i,j)$ (1)
Since the output of convolution represents a locally weighted linear combination of the input, it typically results in reduced sparsity. This means the output generally contains more nonzero elements due to the multiplicative and additive properties of convolution than the input. Therefore, a blurry image has fewer pixels which are very low in brightness. That is, it possesses more nonzero elements compared to the corresponding sharp image.
Now, We define the RFT as a function that processes an image through a series of steps: first, it computes the Fast Fourier Transform (FFT) of the image; then, it applies the ReLU (Rectified Linear Unit) activation function to the FFT results; next, it calculates the inverse FFT of the activated data; finally, it subtracts half of the original image from this result to produce the final output.
@@ 3.1 Sparsity and RFT
[24] show that by applying the aforementioned sequence of operations, faithful information about the blur pattern can be discovered, such as the direction and the level of blur. We observe that the sparsity properties of convolution are preserved even after applying the RFT operations. This observation enables us to penalize the nonzero results from RFT, as these indicate the presence of blur in the image. We apply the RFT operations from the dataset provided by [39] and observe the values of L0 norm to empirically substantiate our observation in Fig. 2. We can observe the change in the L0 norm in blurry images compared to their sharp counterparts when the RFT operation is applied to both.
Figure 2: Decreased Sparsity in Blurry Images after RFT
Here, we can clearly see that the L0 norm of RFT, when applied on blurry images, produces a higher value when compared to applying it on sharp ones.
@@ 3.2 Model and Optimization
Integrating $RFT(I)$ to the standard formulation of image deblurring, we propose a novel objective function as follows:
@@
In this formulation, the first term ensures that the given blurred image and the convolution of the estimated latent image and blur kernels are as similar as possible. The second term applies regularization on the kernel. The third term retains large gradients and discards smaller ones, as discussed in [31, 49]. The parameters $\alpha$, $\beta$, and $\lambda$ serve as penalty coefficients. From Equation (3), we alternately solve for the latent image $I$ and the blur kernel $k$ using coordinate descent, employing the following equations respectively:
@@ 3.3 Blind Deconvolution
In this section, we introduce the methodology for estimating the latent image $I$ and the blur kernel $k$ using the input blurry image $B$. This estimation process is fundamental to the recovery of the latent (sharp) image.
Estimating I: The non-linear $RFT$ function and L0 regularisation make the minimization of Equation (4) challenging in terms of computation. Half-quadratic splitting methods from [47] are used to handle L0 minimization. We split the gradients into $g=(g_{x},g_{y})$ for gradients across x and y-axis, and introduce auxiliary variable $h$ for $RFT(I)$. Rewriting Equation (4), we now have:
$\operatorname*{min}_{I,h,g}||I\otimes k-B||^{2}_{2}+\gamma||\nabla I-g||^{2}_{2}+\beta||RFT(I)-h||^{2}_{2}+\mu||g||_{0}+\lambda||h||_{0}$ (6)
where $\gamma$ and $\beta$ are penalty parameters. Solutions for individual variables ($I,u,$ or $g$) can be calculated when the others are held constant. In order to use Equation (6) in Equation (1) to estimate the latent sharp image, we transform this in the frequency domain and reach a closed-form solution in Equations (10), (11), and (12).
In order to solve for the nonlinear $RFT(I)$, we observe that a linear operator F can be used as a substitute when applied to the I in vector form i.e. for the latent image, FI = RFT(I). To compute the values of F, we use gradient descent with Adam optimizer. Using repeated reconstruction, we iteratively compute $F$ from the previous estimation of $I$. We look at an example image pair from [16] and compare the rate of convergence across different number of steps in Fig. 3.
Figure 3: Convergence of RFT
@@
where $\textbf{T}_{k}$ is a Toeplitz matrix of k, which is multiplied with vectors using FFT B, g and u are denoted in their vector forms, respectively. The solution for Equation (7) is computed as follows:
$L$ $=\operatorname*{min}_{I}||\textbf{T}_{k}\textbf{I}-\textbf{B}||^{2}_{2}+\gamma||\nabla\textbf{I}-\textbf{g}||^{2}_{2}+\beta||\textbf{FI}-\textbf{h}||^{2}_{2}$ (8)
$=(\textbf{T}_{k}\textbf{I}-\textbf{B})^{T}(\textbf{T}_{k}\textbf{I}-\textbf{B})+\gamma(\nabla\textbf{I}-\textbf{g})^{T}(\nabla\textbf{I}-\textbf{g})+$
$\beta(\textbf{FI}-\textbf{h})^{T}(\textbf{FI}-\textbf{h})$
$=\textbf{I}^{T}\textbf{T}_{k}^{T}\textbf{T}_{k}\textbf{I}-\textbf{I}^{T}\textbf{T}_{k}^{T}\textbf{B}-\textbf{B}\textbf{T}_{k}^{T}\textbf{I}+\textbf{B}^{T}\textbf{B}+\gamma\textbf{I}^{T}\nabla^{T}\nabla I-\gamma\textbf{I}^{T}\nabla^{T}g-$
$\gamma\textbf{g}^{T}\nabla\textbf{I}+\gamma\textbf{g}^{T}\textbf{g}+\beta\textbf{I}^{T}\textbf{F}^{T}\textbf{FI}-\beta\textbf{I}^{T}\textbf{F}^{T}\textbf{h}-\beta\textbf{h}^{T}\textbf{FI}+\beta\textbf{h}^{T}\textbf{h}$
Here, $\nabla=(\nabla_{x},\nabla_{y})$ represents the matrix for computing axiswise gradients. Differentiating Equation (8) with respect to $I$ , we get:
$\frac{dL}{dI}$ $=2\textbf{T}_{k}^{T}\textbf{T}_{k}\textbf{I}-\textbf{T}_{k}^{T}\textbf{B}-\textbf{T}_{k}^{T}\textbf{B}+0+2\gamma\nabla^{T}\nabla I-\gamma\nabla^{T}\textbf{g}-\gamma\nabla^{T}\textbf{g}+0+$ (9)
$2\beta\textbf{F}^{T}\textbf{FI}-\beta\textbf{F}^{T}\textbf{h}-\beta\textbf{F}^{T}\textbf{h}$
$=2\textbf{T}_{k}^{T}\textbf{T}_{k}\textbf{I}-2\textbf{T}_{k}^{T}\textbf{B}+2\gamma\nabla^{T}\nabla I-2\gamma\nabla^{T}\textbf{g}+2\beta\textbf{F}^{T}\textbf{FI}-2\beta\textbf{F}^{T}\textbf{h}$
In the case of convergence, $\frac{dL}{dI}=0$. Therefore,
@@
Given I, we can solve for $h$ from:
@@
and for $g$ using element-wise minimization from:
@@
The solution for $h$ can be written as:
@@
Estimating k: Solving Equation (5) becomes a least squares problem when $I$ is fixed. Owing to the demonstrated effectiveness of gradient-based methods in kernel estimation, as detailed in [49, 10, 22], the kernel $k$ is computed from:
@@
FFTs are used to solve Equation (14), using techniques mentioned in [49, 31, 10]. Negative elements of the kernel are set to zero, and the obtained kernel is then normalized.

### v3

# Blind Image Deblurring with FFT-ReLU Sparsity Prior

@@ Abstract
Blind image deblurring is the process of recovering a sharp image from a blurred one without prior knowledge about the blur kernel. It is a small data problem, since the key challenge lies in estimating the unknown degrees of blur from a single image or limited data, instead of learning from large datasets. The solution depends heavily on developing algorithms that effectively model the image degradation process. We introduce a method that leverages a prior which targets the blur kernel to achieve effective deblurring across a wide range of image types. In our extensive empirical analysis, our algorithm achieves results that are competitive with the state-of-the-art blind image deblurring algorithms, and it offers up to two times faster inference, making it a highly efficient solution. [^1] The source code will be available in the authors’ Github profile.
@@ 1 Introduction
Blind image deblurring aims to recover a sharp image $I$ from a blurred image $B$ without knowing the blur kernel $k$. Mathematically, the blurred image $B$ is modeled as:
$B=I\otimes k+\eta$
where $\otimes$ denotes convolution and $\eta$ represents noise. The task is to estimate both $I$ and $k$ from $B$. This is formulated as an optimization problem:
$\min_{I,k}\|B-I\otimes k\|^{2}+\lambda\mathcal{R}(I,k)$
where $\|\cdot\|^{2}$ measures the difference between the observed and predicted images, $\mathcal{R}(I,k)$ is a regularization term to enforce constraints, and $\lambda$ is a regularization parameter. The goal is to recover both $I$ and $k$ by balancing data fidelity and regularization.
Blind image deblurring addresses a fundamental challenge in image processing: restoring clarity from blurred images when the exact nature of the blur is unknown. The algorithmic approach generalizes the image deblurring process, helping data-driven methods perform better.
A common approach in blind image deblurring is the maximum a posteriori (MAP) framework, where the latent image and the blur kernel are alternately optimized. Various heuristics involving the latent image, the kernel, or both have been developed in this framework [3, 10, 20, 23, 32, 35, 48]. This results in new priors that produce effective outcomes. These methods often target specific types of images, as they rely on image properties that aid optimization [32]. Additionally, the computation of some priors can be resource-intensive [33, 4], which can affect the overall efficiency of deblurring techniques. While these approaches demonstrate promising results in their respective domains, their performance can deteriorate when applied to broader image categories, leading to inconsistent outcomes in some cases.
Mao et al. [28] show that applying a specific sequence of operations—removing negative frequencies from a blurry image, reconstructing it in the spatial domain, and subtracting half of the original image—yields an implicit blur kernel representation. This process helps in understanding the blur’s direction and intensity. Building on this insight, we discovered that, when applied to a sharp image, this sequence results in a null kernel, characterized by significantly fewer non-zero elements compared to any blur kernel’s implicit representation.
Our thorough empirical analysis demonstrates that the $L_{0}$ norm obtained after applying this sequence of operations to a blurry image consistently exceeds that of its sharp counterpart. Leveraging this property, we incorporated it into our optimization model, leading to a highly efficient algorithm. The linear nature of our calculation allows our algorithm to be up to two times faster than popular blind image deblurring methods, while still delivering competitive results in terms of PSNR, SSIM, and error ratio.
Figure 1: Results of our blind image deblurring algorithm, compared with other state-of-the-art algorithms. From left to right: 1) Input blurry image, followed by results from 2) Chen et al. [4] 3) Wen et al. [42] 4) Pan et al. [32], 5) Pan et al. [33] and 6) our algorithm.
@@
    We propose a new prior, the ReLU Sparsity prior, which allows us to obtain faithful information about the blur kernel regardless of the nature of the image, while simultaneously applying optimization techniques based on the change in sparsity that occurs due to convolution. Fig. 1 shows an example of deblurred output from our algorithm.
@@
    Our algorithm achieves competitive performance based on the PSNR, SSIM, and error ratio metrics when compared with the state-of-the-art blind image deblurring algorithms [9, 33, 34, 35, 44, 10, 20] while achieving significantly lower inference time on high-resolution images when compared using images from benchmark datasets [18, 22, 38].
In the following sections, we begin by discussing related works pertinent to our study, followed by a detailed explanation of our proposed prior and the empirical evidence supporting the generalizability of the novel property. We then present our experimental results, starting with quantitative analysis and concluding with qualitative evaluations.
2 Related Works
Due to the joint estimation of the blur kernel and the latent sharp image from the input blurry image, single-image deblurring is an under-constrained problem, which is typically formulated in MAP frameworks [27]. Initial advances relied on statistical priors and salient edge detection [9, 44, 48, 35, 10, 20, 23], although the existence of strong edges in latent images is not guaranteed.
Different image priors and likelihood estimations have been proposed to improve the efficiency of MAP frameworks. Such regularizations or assumptions about the blur kernel or the latent image include L0 regularised prior [32, 47], dark channel prior [33], extreme channel prior [49], patch prior [26], local binary pattern prior [51], latent structure prior [2], learned image prior [24], uniform blur [22, 48], non-uniform blur with multiple homographies [14, 30], local maximum difference prior [25], internal patch recurrence [29], Laplacian prior [7], tri-segment intensity prior [50], graph-based image prior [1], constant depth [11, 45], in-plane rotation [37], and forward motion [52]. Fergus et al. [10] use variational Bayesian inference to learn an image gradient prior. Levin et al. [22] show that this method can avoid trivial solutions that naive MAP-based methods can possibly be unable to.
Sparsity priors have been shown to be useful in kernel estimation in MAP frameworks. Krishnan et al. [20] employed a normalized sparsity prior in their MAP framework for kernel estimation. Xu and Jia [44] introduced a two-phase technique for single-image deblurring. First, they estimated the blur kernel using edge selections and ISD optimizations. This was followed by non-blind deconvolution with total variation and an added Gaussian prior [16, 9].
Some comparatively newer approaches designed for domain-specific deblurring, such as low-illumination [13], text [6, 8, 32] and face [31] images utilize inherent statistical characteristics of their domains. Chen et al. [4] propose the local maximum gradient (LMG) prior, which utilizes the diminishing of the maximum value of a local patch during the blurring process. Wen et al. [42] propose the patch-wise minimum pixel (PMP) prior, which is built on the fact that the local minimal pixel would have a higher value in intensity after the blurring process due to the smoothing effect on the image pixels of the process itself. Both these priors use non-linear optimization schemes in MAP frameworks and perform well on both natural and specific images.
3 FFT-ReLU Sparsity Prior
In this section, we discuss the sparsity property of applying RFT operation and develop the FFT-ReLU Sparsity prior to formulate an objective function, in order to estimate the latent sharp image and blur kernel.
3.1 Sparsity and RFT
Sparsity: To describe our work, we begin by discussing the effect of convolution on the sparsity of images. In standard deblurring models, blurry images are denoted as the convolution of the sharp image and a blur kernel. Essentially, the expression can be written as:
$B(x,y)=\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}S(x+i,y+j)\cdot K(i,j)$ (1)
Here, $(x,y)$ denotes the pixel location, and the blur kernel $K$ has a size of $k\times k$. $S$ and $B$ are the sharp image and the blurry image, respectively.
Since the output of convolution essentially represents a locally weighted linear combination of the input, it is essentially less sparse (meaning it has more nonzero elements, due to the multiplicative and additive properties of convolution) than the input. Therefore, a blurry image has fewer pixels which are very low in brightness (i.e. it has more nonzero elements), compared to the corresponding sharp image.
RFT: As introduced by Mao et al. [28], RFT is defined as a function that computes the following for an image:
  1. 1.
    Computes the FFT of the input blurry image
  2. 2.
    Applies the ReLU activation function on it
@@ 3.1 Sparsity and RFT
    Computes the inverse FFT
  4. 4.
    Subtracts half of the image from the result
Formally, we can describe RFT as:
@@
Mao et al. show that faithful information about the blur pattern, such as the direction and the level of blur, can be discovered by applying this sequence of operations. We observe that applying RFT to a sharp image results in fewer non-zero elements than its blurry counterpart, which provides us with the opportunity to penalize the nonzero results from RFT since those indicate the presence of blur.
We apply the RFT operations on the images from various datasets, including one provided by Sun et al. [38], datasets of text images [36], low-light images [41], face images [15] and observe the values of the L0 norm to empirically substantiate our observation in Fig. 2. Since only Sun et al. provided blurry images with their corresponding sharp pairs, we generated blurry images from the provided sharp images for the other datasets by using blur kernels provided by Chen et al. [5] randomly. We can see the change in L0 norm in blurry images from its sharp counterpart when RFT operation is applied to both of them in Fig. 2.
Figure 2: Decreased Sparsity in Blurry Images after RFT, as demonstrated in images from Sun et al. [38], TextOCR dataset, [36] LOL dataset, [41] and LFW dataset [15].
@@ 3.2 Model and Optimization
Adding $RFT(I)$ to standard formulations of image deblurring, we express our objective function as follows:
@@
Here, the first term ensures that the given blurred image and the convolution of the estimated latent image and blur kernels are as similar as possible. The second term applies regularisation on the kernel. The third term retains large gradients and discards smaller ones [32, 48]. $\alpha,\beta$ and $\lambda$ are penalty parameters. From Eq. 3, we solve alternately for the $I$ and $k$ using coordinate descent using following equations respectively:
@@
Our algorithm can be partitioned into blind deconvolution and non-blind deconvolution segments. In blind deconvolution, the computation of the latent image and the blur kernel is done without an exact PSF by using priors or heuristics. The non-blind deconvolution removes the ringing artifacts that can appear from blind deconvolution.
@@ 3.3 Blind Deconvolution
The purpose of blind deconvolution is to estimate the sharp image $I$ and the blur kernel $k$ from the input image $B$.
Estimating I: The non-linear $RFT$ function and L0 regularisation make the minimization of Eq. 4 challenging in terms of computation. Half-quadratic splitting methods [46] are used to handle $L_{0}$ minimization. We split the gradients into $g=(g_{x},g_{y})$ for gradients across x and y-axis, and introduce auxiliary variable $h$ for $RFT(I)$. Rewriting Eq. 4, we now have:
$\operatorname*{min}_{I,h,g}||I\otimes k-B||^{2}_{2}+\gamma||\nabla I-g||^{2}_{2}+\beta||RFT(I)-h||^{2}_{2}\\ +\mu||g||_{0}+\lambda||h||_{0}$ (6)
where $\gamma$ and $\beta$ are penalty parameters. Solutions for individual variables ($I,u,$ or $g$) can be calculated when the others are held constant. In order to use Eq. 6 in Algorithm 1 to estimate the latent sharp image, we transform this in the frequency domain and reach a closed-form solution in Eq. 8, Eq. 9, and Eq. 10.
In order to solve for the nonlinear $RFT(I)$, we observe that a linear operator $F$ can be used as a substitute when applied to the $I$ in vector form i.e. for the latent image, $FI$ = RFT($I$). To compute the values of $F$, we use gradient descent with Adam optimizer [17]. Using repeated reconstruction, we iteratively compute $F$ from the previous estimation of $I$. We look at all images from TextOCR dataset [36], LOL dataset [41], and LFW dataset [15] and observe that $FI$ closely approximates RFT($I$) and converges reasonably within 100 steps in Figure Fig. 3.
@@
where $\textbf{T}_{k}$ is a Toeplitz matrix of k, which is multiplied with vectors using FFT [39]. B, g and u are denoted in their vector forms, respectively. The closed-form solution for $I$ can be derived as follows:
@@
The detailed derivation of this solution is presented in the supplementary material. Given I, we can solve for $h$ from:
@@
and for $g$ from:
@@
using element-wise minimization. The solution for $h$ can be written as:
@@
Estimating k: Eq. 5 is a least squares problem when $I$ is given. Due to the proven accuracy of gradient-based methods for estimating kernels in [48, 9, 23], $k$ is computed from:
@@
FFTs are used for solving Eq. 12, using techniques mentioned in [48, 32, 9]. The negative elements of the kernel are set to zero and the obtained kernel is normalized.
Figure 3: Using gradient descent, FI and RFT(I) converge reasonably within 100 steps.